OpenStudy (anonymous):

how do you solve 16^(-3/4) without a calculator?

OpenStudy (anonymous):

split the powers and then do individually

OpenStudy (anonymous):

what do you mean by split the powers?

OpenStudy (anonymous):

do u got it

OpenStudy (anonymous):

no, im lost where -3 is multiplyed by 6 and then the ^1/4 part

OpenStudy (anonymous):

ok i explain in steps

OpenStudy (theeric):

(-3) means \[\frac{1}{16^3}\] and (1/4) means \[\sqrt[4]{16}\] right? So, since you're doing both to the 16... it can be written as \[\frac{1}{\sqrt[4]{16^3}}\] or \[\frac{1}{\sqrt[4]{16}^3}\]

OpenStudy (anonymous):

that makes sense so far

OpenStudy (theeric):

Now, one looks much more easy to solve than the other, even though they are the same thing.

OpenStudy (anonymous):

the power is -3/4 base is 16 now 16^-3 . 16^1/4=1/(16)^3.\[\sqrt[4]{16}\]

OpenStudy (anonymous):

how does \[\sqrt[4]{16^3}\] turn into 8?

OpenStudy (anonymous):

i know the answer is 1/8

OpenStudy (anonymous):

\[\sqrt[4]{4096}\]

OpenStudy (anonymous):

nvm

OpenStudy (theeric):

I would look at it the other way - just to avoid that huge number \[16^3\].

OpenStudy (theeric):

\[\frac{1}{\sqrt[4]{16}^3}\]

OpenStudy (theeric):

So it's like \[\frac{1}{(\sqrt[4]{16})^3}\]

OpenStudy (theeric):

And what is \[\sqrt[4]{16}\]?

OpenStudy (theeric):

Guess :P

OpenStudy (theeric):

It's a nice number.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

sorry, i was afk for a while

OpenStudy (anonymous):

then 2^3 is 8

OpenStudy (anonymous):

so 1/8

OpenStudy (anonymous):

thanks!

OpenStudy (theeric):

Understood! Thank you! And you're welcome!